OpenStudy (anonymous):

what is the solution of the system? y = 5x + 4 y = 5x - 3

OpenStudy (anonymous):

\[\huge y = 5x + 4 y\] \[\huge y = 5x - 3 \] Are the two equations right?

OpenStudy (anonymous):

Yes they are.

OpenStudy (anonymous):

wait there is no Y after the 4!

OpenStudy (anonymous):

Both are equal to y , So if they both are equal to a specific quantity they themselves must be equal like \[\huge 5x + 4 y = 5x - 3 \]

OpenStudy (anonymous):

I just caught that, sorry!

OpenStudy (anonymous):

what happened?

OpenStudy (anonymous):

is it x then

OpenStudy (anonymous):

the equations are Y = 5x + 4 and Y = 5x - 3

OpenStudy (anonymous):

they want to know how would this be graphed .

OpenStudy (anonymous):

Ok so, you have to graph the equations

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

So I will tell you a general method of graphing any equation 1. You need to substitute "x" the variable on the right hand side with any quantity such 1 ,-1 , 0 , 7 etc 2. Calculate the the respective y values for that x co-ordinate 3. Plot the points on axis and draw a straight line passing through them and you are done

OpenStudy (anonymous):

I will give you a example for the first one , first read what i have written and wether it makes sense

OpenStudy (anonymous):

Wow, its that simple? Thank you!

OpenStudy (anonymous):

Yes its that simple dude!

OpenStudy (anonymous):

So i will give you a example

OpenStudy (anonymous):

y = 5x +4 Substitute any value of x in the equation y = 5(1) +4 Substituted x as one y = 9 1st point :- (1,9) Like that do this for atleast three points to draw a line Although you u need two points to draw any line

OpenStudy (anonymous):

Your graph would look something like this https://www.google.co.in/webhp?sourceid=chrome-instant&ion=1&espv=2&ie=UTF-8#q=y+%3D+5x+%2B+4

OpenStudy (anonymous):

Ohhhhhh okay I see. Thanks! I get it now.

OpenStudy (anonymous):

No problem , you are welcome